# Improving Text Embeddings with Large Language Models

### v2

@@ Abstract
In this paper, we introduce a novel and simple method for obtaining high-quality text embeddings using only synthetic data and less than k training steps. Unlike existing methods that often depend on multi-stage intermediate pre-training with billions of weakly-supervised text pairs, followed by fine-tuning with a few labeled datasets, our method does not require building complex training pipelines or relying on manually collected datasets that are often constrained by task diversity and language coverage. We leverage proprietary LLMs to generate diverse synthetic data for hundreds of thousands of text embedding tasks across nearly languages. We then fine-tune open-source decoder-only LLMs on the synthetic data using standard contrastive loss. Experiments demonstrate that our method achieves strong performance on highly competitive text embedding benchmarks without using any labeled data. Furthermore, when fine-tuned with a mixture of synthetic and labeled data, our model sets new state-of-the-art results on the BEIR and MTEB benchmarks.
@@ 1 Introduction
Previous studies have demonstrated that weighted average of pre-trained word embeddings (Pennington et al., 2014; Arora et al., 2017) is a strong baseline for measuring semantic similarity. However, these methods fail to capture the rich contextual information of natural language. With the advent of pre-trained language models (Devlin et al., 2019), Sentence-BERT (Reimers and Gurevych, 2019) and SimCSE (Gao et al., 2021) have been proposed to learn text embeddings by fine-tuning BERT on natural language inference (NLI) datasets. To further enhance the performance and robustness of text embeddings, state-of-the-art methods like E5 (Wang et al., 2022b) and BGE (Xiao et al., 2023) employ a more complex multi-stage training paradigm that first pre-trains on billions of weakly-supervised text pairs, and then fine-tunes on several labeled datasets.
@@
In this paper, we propose a novel method for text embeddings that leverages LLMs to overcome the limitations of existing approaches. We use proprietary LLMs to generate synthetic data for a diverse range of text embedding tasks in languages, covering hundreds of thousands of embedding tasks. Specifically, we use a two-step prompting strategy that first prompts the LLMs to brainstorm a pool of candidate tasks, and then prompts the LLMs to generate data conditioned on a given task from the pool. To cover various application scenarios, we design multiple prompt templates for each task type and combine the generated data from different templates to boost diversity. For the text embedding models, we opt for fine-tuning powerful open-source LLMs rather than small BERT-style models. Since LLMs such as Mistral (Jiang et al., 2023) have been extensively pre-trained on web-scale data, contrastive pre-training offers little additional benefit.
@@ 2 Related Work
Text Embeddings are continuous low-dimensional representations of text and have been extensively applied to various downstream tasks such as information retrieval, question answering, and retrieval-augmented generation (RAG). Early work on text embeddings includes latent semantic indexing (Deerwester et al., 1990) and weighted average of word embeddings (Mikolov et al., 2013). More recent methods exploit supervision from natural language inference (Bowman et al., 2015) and labeled query-document pairs, such as the MS-MARCO passage ranking dataset (Campos et al., 2016), to train text embeddings (Reimers and Gurevych, 2019; Conneau et al., 2017; Gao et al., 2021). However, labeled data are often limited in terms of task diversity and language coverage. To address this challenge, methods like Contriever (Izacard et al., 2021), OpenAI Embeddings (Neelakantan et al., 2022), E5 (Wang et al., 2022b), and BGE (Xiao et al., 2023) adopt a multi-stage training paradigm. They first pre-train on large-scale weakly-supervised text pairs using contrastive loss and then fine-tune on small-scale but high-quality datasets. In this paper, we demonstrate that it is possible to obtain state-of-the-art text embeddings with single-stage training.
@@
Another related line of work focuses on knowledge distillation from black-box LLMs by training on synthetic data generated from them. DINO (Schick and Schütze, 2021) generates synthetic text pairs for semantic textual similarity. Unnatural Instructions (Honovich et al., 2022) is a synthetic instruction following dataset by prompting existing LLMs. Orca (Mukherjee et al., 2023) and Phi (Gunasekar et al., 2023) propose to train better small language models by using high-quality synthetic data from GPT-3.5/4 (OpenAI, 2023).
@@ 3.1 Synthetic Data Generation
To generate diverse synthetic data, we propose a simple taxonomy that categorizes embedding tasks into several groups, and then apply different prompt templates to each group.
Asymmetric Tasks This category comprises tasks where the query and document are semantically related but are not paraphrases of each other. Depending on the length of the query and document, we further divide asymmetric tasks into four subgroups: short-long match, long-short match, short-short match, and long-long match. For instance, short-long match tasks involve a short query and a long document, which is a typical scenario in commercial search engines. For each subgroup, we design a two-step prompt template that first prompts LLMs brainstorm a list of tasks, and then generates a concrete example conditioned on the task definition. In Figure 1, we show an example prompt for the short-long match subgroup. The outputs from GPT-4 are mostly coherent and of high quality. In our preliminary experiments, we also attempted to generate the task definition and query-document pairs using a single prompt, but the data diversity was not as satisfactory as the proposed two-step approach.
Symmetric Tasks Symmetric tasks involve queries and documents that have similar semantic meanings but different surface forms. We examine two application scenarios: monolingual semantic textual similarity (STS) and bitext retrieval. We design two distinct prompt templates for each scenario, tailored to their specific objectives. Since the task definition is straightforward, we omit the brainstorming step for symmetric tasks.
@@ 4.1 Statistics of the Synthetic Data
Figure 2 presents the statistics of our generated synthetic data. We manage to generate k examples with k unique instructions using Azure OpenAI Service 111https://oai.azure.com/, among which are generated by GPT-35-Turbo and others are generated by GPT-4. The total token consumption is about M. The predominant language is English, with coverage extending to a total of languages. For the bottom low-resource languages, there are about k examples per language on average.
4.2 Model Fine-tuning and Evaluation
4.3 Main Results
@@
| Model | BEIR Retrieval (15 datasets) | MTEB Average (56 datasets) |
| OpenAI Ada-002 | 49.3 | 61.0 |
| Cohere-embed-english-v3.0 | 55.0 | 64.5 |
| voyage-lite-01-instruct | 55.6 | 64.5 |
| UAE-Large-V1 (Li and Li, 2023) | 54.7 | 64.6 |
| E5 + full data | 56.9 | 66.6 |
In Table 2, we also present a comparison with several commercial text embedding models. However, due to the lack of transparency and documentation about these models, a fair comparison is not feasible. We focus especially on the retrieval performance on the BEIR benchmark, since RAG is an emerging technique to enhance LLM with external knowledge and proprietary data. As Table 2 shows, our model outperforms the current commercial models by a significant margin.
4.4 Multilingual Retrieval
@@
To assess the multilingual capabilities of our model, we conduct an evaluation on the MIRACL dataset (Zhang et al., 2023b), which comprises human-annotated queries and relevance judgments across languages. As shown in Table 3, our model surpasses mE5 on high-resource languages, notably on English. Nevertheless, for low-resource languages, our model remains suboptimal compared to mE5. We attribute this to the fact that Mistral-7B is predominantly pre-trained on English data, and we anticipate that future multilingual LLMs will leverage our method to bridge this gap.
5 Analysis
5.1 Is Contrastive Pre-training Necessary?
5.2 Extending to Long Text Embeddings
5.3 Analysis of Training Hyperparameters
@@
Table 4 presents the results under different configurations. We notice that the Mistral-7B initialization holds an advantage over LLaMA-2 7B, in line with the findings from Mistral-7B technical report (Jiang et al., 2023). The choice of pooling types and LoRA ranks does not affect the overall performance substantially, hence we adhere to the default setting despite the marginal superiority of LoRA rank . On the other hand, the way of adding instructions has a considerable impact on the performance. We conjecture that natural language instructions better inform the model regarding the embedding task at hand, and thus enable the model to generate more discriminative embeddings. Our framework also provides a way to customize the behavior of text embeddings through instructions without the need to fine-tune the model or re-built document index.
@@ 6 Conclusion
For future work, we aim to further improve the multilingual performance of our model and explore the possibility of using open-source LLMs to generate synthetic data. We also intend to investigate ways to improve the inference efficiency and lower the storage cost for LLM based text embeddings.
@@ Appendix A Implementation Details
Baseline Models For results with mE5 and mE5, we use the public checkpoints available at https://huggingface.co/intfloat/multilingual-e5-base and https://huggingface.co/intfloat/multilingual-e5-large respectively. For experiments in Table 4, we follow the SGPT [Muennighoff, 2022] paper for the implementation of weighted mean pooling. For the “w/ task type prefix” setting, we prepend “classify: ” for the long-short matching subgroup, and “query: ” for other asymmetric tasks. No prefix is added for symmetric tasks.
Training Data For the “E5 + full data” setting, our training data comprises generated synthetic data, ELI5 [Fan et al., 2019](sample ratio ), HotpotQA [Yang et al., 2018], FEVER [Thorne et al., 2018], MIRACL [Zhang et al., 2023b], MSMARCO passage ranking (sample ratio ) and document ranking (sample ratio ) [Campos et al., 2016], NQ [Karpukhin et al., 2020], NLI [Gao et al., 2021], SQuAD [Karpukhin et al., 2020], TriviaQA [Karpukhin et al., 2020], Quora Duplicate Questions [DataCanary et al., 2017](sample ratio ), MrTyDi [Zhang et al., 2021], DuReader [Qiu et al., 2022], and T2Ranking [Xie et al., 2023](sample ratio ) datasets. We only include the training set of each dataset. For the datasets without hard negatives, we use mE5 to mine top hard negatives. After sampling, we obtain approximately million examples. The entire training process takes fewer than k steps to complete.
Hyperparameters for Fine-tuning When fine-tuning Mistral-7b, the batch size is set to and the learning rate is with step warmup and linear decay. The weight decay is . We add hard negative for each query-document pair. The fine-tuning process takes roughly hours on V100 GPUs with a maximum sequence length . We add LoRA adapters to all linear layers, resulting in a total of M trainable parameters. Our implementation is based on the HuggingFace PEFT library at https://github.com/huggingface/peft.
The model and dataset release information is available at https://github.com/microsoft/unilm/tree/master/e5.
Appendix B Test Set Contamination Analysis
@@ Appendix C Prompts for Synthetic Data Generation
For asymmetric tasks, we list the four prompt templates in Table 7, 8, 9, and 10. For symmetric tasks, the prompts templates are available in Table 11 and 12. To generate multilingual data, we sample the value of “{language}” from the language list of XLM-R [Conneau et al., 2020] with higher probability for high-resource languages. When prompting GPT-4/3.5, we set the temperature to and the top- to , which is higher than the default setting to encourage more diversity.

### v3

@@ Abstract
In this paper, we introduce a novel and simple method for obtaining high-quality text embeddings using only synthetic data and less than k training steps. Unlike existing methods that often depend on multi-stage intermediate pre-training with billions of weakly-supervised text pairs, followed by fine-tuning with a few labeled datasets, our method does not require building complex training pipelines or relying on manually collected datasets that are often constrained by task diversity and language coverage. We leverage proprietary LLMs to generate diverse synthetic data for hundreds of thousands of text embedding tasks across languages. We then fine-tune open-source decoder-only LLMs on the synthetic data using standard contrastive loss. Experiments demonstrate that our method achieves strong performance on highly competitive text embedding benchmarks without using any labeled data. Furthermore, when fine-tuned with a mixture of synthetic and labeled data, our model sets new state-of-the-art results on the BEIR and MTEB benchmarks.
Improving Text Embeddings with Large Language Models
Liang Wang, Nan Yang, Xiaolong Huang, Linjun Yang, Rangan Majumder, Furu Wei Microsoft Corporation {wangliang,nanya,xiaolhu,yang.linjun,ranganm,fuwei}@microsoft.com
@@ 1 Introduction
Previous studies have demonstrated that weighted average of pre-trained word embeddings (Pennington et al., 2014; Arora et al., 2017) is a strong baseline for measuring semantic similarity. However, these methods fail to capture the rich contextual information of natural language. With the advent of pre-trained language models (Devlin et al., 2019), Sentence-BERT (Reimers and Gurevych, 2019) and SimCSE (Gao et al., 2021) have been proposed to learn text embeddings by fine-tuning BERT on natural language inference (NLI) datasets. To further enhance the performance and robustness of text embeddings, state-of-the-art methods like E5 (Wang et al., 2022b) and BGE (Xiao et al., 2023) employ a more complex multi-stage training paradigm that first pre-trains on billions of weakly-supervised text pairs, and then fine-tunes on several high-quality labeled datasets.
@@
In this paper, we propose a novel method for text embeddings that leverages LLMs to overcome the limitations of existing approaches. We use proprietary LLMs to generate synthetic data for a diverse range of text embedding tasks in languages, covering hundreds of thousands of embedding tasks. Specifically, we use a two-step prompting strategy that first prompts the LLMs to brainstorm a pool of candidate tasks, and then prompts the LLMs to generate data conditioned on a given task from the pool. To cover various application scenarios, we design multiple prompt templates for each task type and combine the generated data from different templates to boost diversity. For the text embedding models, we opt for fine-tuning powerful open-source LLMs rather than small BERT-style models. Since LLMs such as Mistral (Jiang et al., 2023) have been extensively pre-trained on web-scale data, contrastive pre-training that proves to be important for BERT models (Wang et al., 2022b) offers little additional benefit.
@@ 2 Related Work
Text Embeddings
are continuous low-dimensional representations of text and
have been extensively applied to various downstream tasks such as information retrieval, question answering,
and retrieval-augmented generation (RAG).
Early work on text embeddings includes latent semantic indexing (Deerwester et al., 1990)
and weighted average of word embeddings (Mikolov et al., 2013).
More recent methods exploit supervision from natural language inference (Bowman et al., 2015) and
labeled query-document pairs, such as the MS-MARCO passage ranking dataset (Campos et al., 2016),
to train text embeddings (Reimers and Gurevych, 2019; Conneau et al., 2017; Gao et al., 2021).
However,
labeled data are often limited in terms of task diversity and language coverage.
To address this challenge,
methods like Contriever (Izacard et al., 2021), OpenAI Embeddings (Neelakantan et al., 2022), E5 (Wang et al., 2022b), and BGE (Xiao et al., 2023)
adopt a multi-stage training paradigm.
They first pre-train on large-scale weakly-supervised text pairs using contrastive loss
and then fine-tune on small-scale but high-quality datasets.
In this paper,
we demonstrate that it is possible to obtain state-of-the-art text embeddings with single-stage training.
@@
Another related line of work focuses on knowledge distillation from black-box LLMs
by training on synthetic data generated from them.
DINO (Schick and Schütze, 2021) generates synthetic text pairs for semantic textual similarity.
Unnatural Instructions (Honovich et al., 2022) is a synthetic instruction following dataset
by prompting existing LLMs.
Orca (Mukherjee et al., 2023) and Phi (Gunasekar et al., 2023) propose to train better small language models
by using high-quality synthetic data from GPT-3.5/4 (OpenAI, 2023).
@@ 3.1 Synthetic Data Generation
To generate diverse synthetic data,
we propose a simple taxonomy that categorizes embedding tasks into several groups,
and then apply different prompt templates to each group.
Asymmetric Tasks
This category comprises tasks where
the query and document are semantically related but are not paraphrases of each other.
Depending on the length of the query and document,
we further divide asymmetric tasks into four subgroups:
short-long match, long-short match, short-short match, and long-long match.
For instance,
short-long match tasks involve a short query and a long document,
which is a typical scenario in commercial search engines.
For each subgroup,
we design a two-step prompt template that first prompts LLMs brainstorm a list of tasks,
and then generates a concrete example conditioned on the task definition.
In Figure 1,
we show an example prompt for the short-long match subgroup.
The full output is available in Table 16.
The outputs from GPT-4 are mostly coherent and of high quality.
In our preliminary experiments,
we also attempted to generate the task definition and query-document pairs using a single prompt,
but the data diversity was not as satisfactory as the proposed two-step approach.
Symmetric Tasks
Symmetric tasks involve queries and documents that have similar semantic meanings but different surface forms.
We examine two application scenarios:
monolingual semantic textual similarity (STS) and bitext retrieval.
We design two distinct prompt templates for each scenario,
tailored to their specific objectives.
Since the task definition is straightforward,
we omit the brainstorming step for symmetric tasks.
@@ 4.1 Statistics of the Synthetic Data
Figure 2 presents the statistics of our generated synthetic data. We manage to generate k examples with k unique instructions using Azure OpenAI Service 111https://oai.azure.com/, among which are generated by GPT-35-Turbo and others are generated by GPT-4. The total token consumption is about M. The predominant language is English, with coverage extending to a total of languages. For the bottom low-resource languages, there are about k examples per language on average. Please see Table 16 in the appendix for examples of synthetic data.
4.2 Model Fine-tuning and Evaluation
@@
4.3 Main Results
| Model | BEIR | MTEB |
| OpenAI text-embedding-3-large | 55.4 | 64.6 |
| Cohere-embed-english-v3.0 | 55.0 | 64.5 |
| voyage-lite-01-instruct | 55.6 | 64.5 |
| UAE-Large-V1 | 54.7 | 64.6 |
| E5 + full data | 56.9 | 66.6 |
In Table 3, we also present a comparison with several commercial text embedding models. However, due to the lack of transparency and documentation about these models, a fair comparison is not feasible. We focus especially on the retrieval performance on the BEIR benchmark, since retrieval-augmented generation is an emerging technique to enhance LLM with external knowledge and proprietary data. As Table 3 shows, our model outperforms the current commercial models by a significant margin.
4.4 Multilingual Retrieval
To assess the multilingual capabilities of our model, we conduct an evaluation on the MIRACL dataset (Zhang et al., 2023b), which comprises human-annotated queries and relevance judgments across languages. The validation set contains labels for languages. As shown in Table 2, our model surpasses mE5 on high-resource languages, notably on English. Nevertheless, for low-resource languages, our model remains suboptimal compared to mE5. We attribute this to the fact that Mistral-7B is predominantly pre-trained on English data, and we anticipate that future multilingual LLMs will leverage our method to bridge this gap.
| mContriever | 93.7 | 37.7 | ||||
| LaBSE | 98.8 | 81.1 | ||||
| mE5 | 98.1 | 68.1 | ||||
| mE5 | 98.6 | 75.7 | ||||
| E5 | 98.9 | 70.1 |
To evaluate our model’s cross-lingual retrieval capability, we report Bitext mining results in Table 4. For baselines including mContriever (Izacard et al., 2021), LaBSE (Feng et al., 2022), and mE5 (Wang et al., 2024), we evaluate the results using publicly available checkpoints. Our observations indicate that, similar to the MIRACL retrieval, E5 excels in bitext mining for high-resource languages only.
@@ 4.4 Multilingual Retrieval
5 Analysis
5.1 Is Contrastive Pre-training Necessary?
5.2 Extending to Long Text Embeddings
5.3 Analysis of Training Hyperparameters
Table 5 presents the results under different configurations. We notice that the Mistral-7B initialization holds an advantage over LLaMA-2 7B, in line with the findings from Mistral-7B technical report (Jiang et al., 2023). The choice of pooling types and LoRA ranks does not affect the overall performance substantially, hence we adhere to the default setting despite the marginal superiority of LoRA rank . On the other hand, the way of adding instructions has a considerable impact on the performance. We conjecture that natural language instructions better inform the model regarding the embedding task at hand, and thus enable the model to generate more discriminative embeddings. Our framework also provides a way to customize the behavior of text embeddings through instructions without the need to fine-tune the model or re-build document index.
@@ 6 Conclusion
For future work, we aim to further improve the multilingual performance of our model and explore the possibility of using open-source LLMs to generate synthetic data.
Limitations
In comparison to the mainstream BERT-style encoders, the employment of LLMs, such as Mistral-7B, for text embeddings results in a significantly increased inference cost. The development of more advanced GPUs and better kernel implementations may enhance the efficiency of the inference process. With regards to storage cost, our model is comparatively more expensive, with embeddings of dimensions. Early successes in reducing embedding dimensions while maintaining competitive performance have been demonstrated through techniques such as Matryoshka representation learning (Kusupati et al., 2022).
For synthetic data generation, we rely on manual prompt engineering to elicit high-quality outputs from proprietary LLMs. Automatic prompt optimization presents a promising avenue for improving the quality of synthetic data.
Acknowledgements
We would like to thank anonymous reviewers for their valuable comments, and ACL 2024 and ACL Rolling Review organizers for their efforts. Opinions expressed in this paper are solely those of the authors and do not represent the views of their employers.
@@ Appendix A Implementation Details
Baseline Models
For results with mE5 and mE5,
we use the public checkpoints available
at https://huggingface.co/intfloat/multilingual-e5-base and https://huggingface.co/intfloat/multilingual-e5-large respectively.
For experiments in Table 5,
we follow the SGPT (Muennighoff, 2022) paper for the implementation of weighted mean pooling.
For the “w/ task type prefix” setting,
we prepend “classify: ” for the long-short matching subgroup,
and “query: ” for other asymmetric tasks.
No prefix is added for symmetric tasks.
Training Data
For the “E5 + full data” setting,
our training data comprises generated synthetic data, ELI5 (Fan et al., 2019)(sample ratio ), HotpotQA (Yang et al., 2018), FEVER (Thorne et al., 2018),
MIRACL (Zhang et al., 2023b),
MSMARCO passage ranking (sample ratio ) and document ranking (sample ratio ) (Campos et al., 2016),
NQ (Karpukhin et al., 2020),
NLI (Gao et al., 2021), SQuAD (Karpukhin et al., 2020), TriviaQA (Karpukhin et al., 2020),
Quora Duplicate Questions (DataCanary et al., 2017)(sample ratio ), MrTyDi (Zhang et al., 2021),
DuReader (Qiu et al., 2022), and T2Ranking (Xie et al., 2023)(sample ratio ) datasets.
We only include the training set of each dataset.
For the datasets without hard negatives,
we use mE5 to mine top hard negatives.
After sampling,
we obtain approximately million examples.
The entire training process takes fewer than k steps to complete.
Hyperparameters for Fine-tuning
When fine-tuning Mistral-7b 222https://huggingface.co/mistralai/Mistral-7B-v0.1,
the batch size is set to and the learning rate is with step warmup and linear decay.
The weight decay is .
We add hard negative for each query-document pair.
The fine-tuning process takes roughly hours on V100 GPUs with a maximum sequence length .
We add LoRA adapters to all linear layers,
resulting in a total of M trainable parameters.
Our implementation is based on the HuggingFace PEFT library at https://github.com/huggingface/peft.
Artifacts The model and dataset release information is available at https://github.com/microsoft/unilm/tree/master/e5. We release our trained models and evaluation scripts to facilitate reproducibility and further research.
Appendix B Test Set Contamination Analysis
@@ Appendix C Prompts for Synthetic Data Generation
For asymmetric tasks, we list the four prompt templates in Table 8, 9, 10, and 11. For symmetric tasks, the prompts templates are available in Table 12 and 13. To generate multilingual data, we sample the value of “{language}” from the language list of XLM-R (Conneau et al., 2020) with higher probability for high-resource languages. When prompting GPT-4/3.5, we set the sampling temperature to and the top- hyperparameter to , which is higher than the default setting to encourage more diversity.